# A Digital Twin of Evaporative Thermo-Fluidic Process in Fixation Unit of DoD Inkjet Printers

### v1

@@ 1.2 Challenges in current fixation process
Although temperature sensors monitor the conveyor, direct measurement of the residual temperature and moisture content in the paper sheets during fixation is not possible because it is impossible for the sensors to be directly in contact with the paper sheets. Furthermore, the high throughput of commercial printing requires fast movement of the paper sheets during fixation, making alternative sensing such as thermal imaging infeasible. However, these spatio-temporal thermo-fluidic effects on the paper sheets are crucial to assess the thermo-mechanical properties of the printed paper sheets as well as the integration of ink to the medium, two key performance indices for the print quality. The state-of-the-art practice in the printing industry largely focuses on design improvements when it comes to monitor and compensate for the thermo-fluidic effects on the paper sheets during fixation, which currently poses as a major quality-limiting factor in DoD inkjet printing. A system-theoretic perspective to model and control thermo-fluidic effects is sorely missing in DoD inkjet technology (9465747). A limited number of efforts for model-based feedforward compensation for these thermo-fluidic effects have been made recently. However, they have been futile because of the absence of accurate spatio-temporal estimation of these effects over the course of fixation. Since Partial Differential Equations (PDEs) are used to model these thermo-fluidic phenomena (pletcher2012computational), the main computational challenge arises from their infinite dimensional nature. Standard synthesis approaches based on early lumping (SHANG2000533; 10.1007/978-3-0348-8849-3_12) and late lumping suffer from the same issue: an inaccurate assessment of robustness and stability certificates, which makes them unreliable for commercial implementation. In das_2019CDC; 2024arXiv241101793B, a state estimation strategy on infinite dimensions is provided which circumvents the issues related to lumping by converting the PDEs to an equivalent set of Partial Integral Equations (PIEs). This synthesis strategy does not involve any approximation and can be solved by convex optimization. However, it is important to note that, beyond a proof of concept, the PIE framework has never been incorporated into industrial products.
@@ 1.3 Contribution
This work develops a modular, physics-based digital twin for simulation, rapid-prototyping and in-operation monitoring the thermo-fluidic dynamics of paper sheets during the fixation process, addressing high throughput in commercial printing. By eliminating the need for additional physical sensors, this digital twin provides a virtual model for experimentation, prediction, and diagnostics, optmization of system responsiveness and reducing reliance on physical testing. To overcome the limitations of existing methods, which struggle with accurate predictions due to the inability to precisely model spatio-temporal thermo-fluidic effects, this study employs Partial Integral Equations (PIEs), a class of equations derived from Partial Differential Equations (PDEs) through a variable transformation. Unlike traditional approaches, PIEs do not require boundary conditions, and their formulation allows the use of Linear Matrix Inequality methods that are widely-used for Ordinary Differential Equations (ODEs). Furthermore, the digital twin incorporates an $\mathcal{H}_{\infty}$-optimal state estimation technique to monitor infinite dimensional states while accounting for the sparse sensor data and unknown external perturbations. By integrating a state estimator with a modular physical model, validated with machine data, this work provides the very first evidence that the PIE framework is implementable in an industrial setting. In this way, the paper offers a pathway to exploit new computational techniques in the realm of distributed parameter systems to significantly improve the quality of DoD inkjet printing processes by more accurately modeling and estimation of the thermo-mechanical interactions during fixation.
@@ 2 Preliminaries
Due to confidentiality reasons, in this paper, all the plots are normalized by subtracting the data points with a fixed nominal value.
In the paper, $\kappa(\mathrm{W/m.K})$ is defined as the thermal conductivity of the material, $\rho(\mathrm{kg/m^{3}})$ is the density of the material, $\chi(\mathrm{J/kg.K})$ is the specific heat capacity, $h(\mathrm{W/m^{2}.K})$ is the heat transfer coefficient, $g(\mathrm{m/s})$ is the moisture transfer coefficient, and $\delta(\mathrm{m^{2}/s})$ is the moisture diffusion coefficient. The space of functions $f:A\rightarrow B$ is denoted by $(B)^{A}$, i.e., $f\in(B)^{A}$. The array $\begin{bmatrix}a\\ b\end{bmatrix}$ is often written as $\text{col}(a,b)$. $\partial^{q}_{s}f$ denotes the qth-order partial derivative of the function $f$ with respect to $s$, i.e. $\frac{\partial^{q}f}{\partial s^{q}}$.
Given bounded matrix $P\in\mathbb{R}^{n\times m}$, and bounded matrix-valued polynomials $\mathbf{Q_{1}}:[\alpha,\beta]\to\mathbb{R}^{n\times p}$, $\mathbf{Q_{2}}:[\alpha,\beta]\to\mathbb{R}^{q\times m}$, $\mathbf{R_{0}}$$\mathbf{R_{1}}$, and $\mathbf{R_{2}}:[\alpha,\beta]\to\mathbb{R}^{q\times p}$, a PI operator denoted by $\mathfrak{\P}\hskip-3.44444pt\left[\footnotesize\begin{array}[]{c|c}P&\mathbf{Q_{1}}\\ \hline\cr\mathbf{Q_{2}}&\{\mathbf{R_{i}}\}\end{array}\right]$ is defined for all $s\in[\alpha,\beta]$, $u\in\mathbb{R}^{m}$ and $\mathbf{v}:[\alpha,\beta]\to\mathbb{R}^{p}$ according to
@@ 3.1 Graph theoretic representation
$\mathcal{G}=(\mathbb{N},A,\mathbb{E}),$
@@ 3.1.1 Node Structure:
$\mathcal{N}_{i}=(\mathbb{X}_{i},\mathbb{X}_{i}^{\mathrm{bc}},\mathfrak{S}_{i},\mathfrak{P}_{i},\mathfrak{P}_{i}^{\mathrm{bc}},\mathfrak{D}_{i}),$
@@
$z_{i}$ Regulated output(C) $\mathbb{T}\rightarrow\mathbb{R}^{n_{z,i}}$ Performance index
@@ 3.1.2 Edge Structure:
$\mathcal{E}_{i,j}=(\mathbb{X}_{i,j}^{I},\mathfrak{C}_{i,j}^{I},\mathfrak{E}_{i,j}),$
@@ 3.2 Graph-Theoretic Model of the Fixation Unit
Figure 2: The graph theoretical representation of the paper-conveyor system. X,X,X represent the paper, conveyor and the air layers respectively. $w_{1}(t)$ is the unknown disturbance, $d_{1}(t),d_{2}(t)$ are the known disturbance acting at the boundaries. $\Delta$ is the non-linearity of the system.
The paper sheet is modeled as a non-homogeneous composite with layers $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$, while node $\mathcal{N}_{5}$ represents the conveyor, where only temperature is defined. The dynamics at each spatially distributed node are governed by Fick’s law, subject to interface and external boundary conditions. Thermal contact at interfaces may be either perfect or imperfect. Specifically, the interaction between the paper and the conveyor is modeled with imperfect thermal contact due to energy loss, whereas the thermal contact between the paper layers is assumed to be perfect. External boundaries may experience disturbances that affect the system.
@@
The moisture content in printed paper is lost through evaporation. The mass flux of the evaporation is a function of paper temperature and moisture and is given by the static non-linear function $\Delta_{1}(p(t))$. The enthalpy of evaporation is also a function of the paper temperature and is given by the non-linear function $\Delta_{2}(p(t))$. This is represented by the Linear Fractional Representation (LFR). The exact $\Delta_{1}(p(t))$ and $\Delta_{2}(p(t))$ function are not mentioned as it is an intellectual property of Canon Production Printing.
Figure 3: Equivalence between the developed model and PIEs
@@
Here, $\mathscr{T},\mathscr{T}_{w},\mathscr{T}_{d}$, $\mathscr{A},\mathscr{B}_{11},\mathscr{B}_{12},\mathscr{B}_{q}$, $\mathscr{C}_{1},\mathscr{D}_{11},\mathscr{D}_{12},\mathscr{D}_{1q}$, $\mathscr{C}_{2},\\ \mathscr{D}_{21},\mathscr{D}_{22},\mathscr{D}_{2q}$, $\mathscr{C}_{p},\mathscr{D}_{p1},\mathscr{D}_{p2},\mathscr{D}_{pq}$ are specific 4-PI operators of appropriate dimensions.
@@
The process of constructing the PIEs from the graph-theoretic representation consists of three steps and is depicted in Fig. 3
@@
For a well-defined PDE, there must be a sufficient number of boundary conditions. Without this, PDE models do not admit a PIE representation. Whether a set of boundary condition is admissible for PIE conversion or not can be easily tested by the invertibility of the matrix $B_{T}$ as given by Theorem 6.2 (phdthesis) where
@@
Since the $B_{T}$ is invertible, using Theorm 6.3 from (phdthesis), the PIE operators $\mathscr{T},\mathscr{T}_{w},\mathscr{T}_{d}$, $\mathscr{A},\mathscr{B}_{11},\mathscr{B}_{12},\mathscr{B}_{q}$, $\mathscr{C}_{1},\mathscr{D}_{11},\mathscr{D}_{12},\mathscr{D}_{1q}$, $\mathscr{C}_{2},\mathscr{D}_{21},\mathscr{D}_{22},\mathscr{D}_{2q}$, $\mathscr{C}_{p},\mathscr{D}_{p1},\mathscr{D}_{p2},\mathscr{D}_{pq}$ are obatined.
@@ 3.3 Implementation and simulation of the model
Based on graph-theoretic definitions, all signals associated with each node are known. The digital twin is implemented in an object-oriented framework using MATLAB and built on top of PIETOOLS (shivakumar2020pietoolsmatlabtoolboxmanipulation). Each node is represented as a class encapsulating its properties to reduce complexity. The layer class defines each paper layer with parameters such as diffusion, transport, reaction, and transfer coefficients, spatial domain, boundary/initial conditions, and boundary inputs. The spatialconveyor class defines spatial conveyor nodes, storing parameters such as diffusion coefficient, domain size, and initial conditions. This modular design allows flexible definition of any number of states per node in arbitrary order, while ensuring correct interconnections through the $E_{i,j}$ matrix between adjacent nodes. The thermo-fluidic process is simulated using the equivalent PIE formulation from (11), solved via PIE-Galerkin Projection (PGP) (PEET2024115673).

### v2

@@ 1.2 Challenges in current fixation process
Although temperature sensors monitor the conveyor, direct measurement of the residual temperature and moisture content in the paper sheets during fixation is not possible because it is impossible for the sensors to be directly in contact with the paper sheets. Furthermore, the high throughput of commercial printing requires fast movement of the paper sheets during fixation, making alternative sensing such as thermal imaging infeasible. However, these spatio-temporal thermo-fluidic effects on the paper sheets are crucial to assess the thermo-mechanical properties of the printed paper sheets as well as the integration of ink to the medium, two key performance indices for the print quality. The state-of-the-art practice in the printing industry largely focuses on design improvements when it comes to monitor and compensate for the thermo-fluidic effects on the paper sheets during fixation, which currently poses as a major quality-limiting factor in DoD inkjet printing. A system-theoretic perspective to model and control thermo-fluidic effects is sorely missing in DoD inkjet technology (9465747). A limited number of efforts for model-based feedforward compensation for these thermo-fluidic effects have been made recently. However, they have been futile because of the absence of accurate spatio-temporal estimation of these effects over the course of fixation. Since Partial Differential Equations (PDEs) are used to model these thermo-fluidic phenomena (pletcher2012computational), the main computational challenge arises from their infinite dimensional nature. Standard synthesis approaches based on early lumping (SHANG2000533; 10.1007/978-3-0348-8849-3_12) and late lumping suffer from the same issue: an inaccurate assessment of robustness and stability certificates, which makes them unreliable for commercial implementation. In 9029595; 2024arXiv241101793B, a state estimation strategy on infinite dimensions is provided which circumvents the issues related to lumping by converting the PDEs to an equivalent set of Partial Integral Equations (PIEs). This synthesis strategy does not involve any approximation and can be solved by convex optimization (also see PEET2019132, sachin). However, it is important to note that, beyond a proof of concept, the PIE framework has never been incorporated into industrial products.
@@ 1.3 Contribution
The fixation unit in commercial DoD inkjet printers must be tuned and evaluated across a wide range of operational scenarios, yet performing repeated physical experiments is both time‑consuming and costly. A physics‑based virtual replica of the fixation process enables the prediction of drying behaviour and the identification of optimal fixation parameters, thereby allowing system-level optimisation in a cost-efficient manner without the need for repeated physical experimentation. A digital twin is such a software-based replica that represents the underlying thermo-fluidic processes in the fixation unit with a multitude of functionalities that are relevant and are customizable based on application-specific requirements.
This work develops a modular, physics‑based digital twin for simulation, rapid prototyping, and in‑operation monitoring of the thermo‑fluidic dynamics of paper sheets during the fixation process. To support this modularity, the physical system is represented using a graph-theoretic structure in which each subsystem is modelled as a node governed by its own PDEs. The purpose of this representation is to provide a scalable organisational framework that cleanly encodes interconnections, boundary interactions, and heterogeneous material layers. This is particularly relevant in industrial printing, where paper may consist of multiple composite layers with different thermo‑fluidic characteristics. The graph-based formulation, therefore, enables straightforward adaptation to different media and fixation-unit configurations, ensuring that the digital twin remains extensible and industry relevant.
To overcome the limitations of existing methods, which struggle with accurate predictions due to the inability to precisely model spatio-temporal thermo-fluidic effects, this study employs PIEs, a class of equations derived from PDEs through a variable transformation. Unlike traditional approaches, PIEs do not require boundary conditions, and their formulation allows the use of Linear Matrix Inequality methods that are widely-used for Ordinary Differential Equations. Furthermore, the digital twin incorporates an $\mathcal{H}_{\infty}$-optimal state estimation technique to monitor infinite dimensional states while accounting for the sparse sensor data and unknown external perturbations. By integrating a state estimator with a modular physical model, validated with machine data, this work provides the very first evidence that the PIE framework is implementable in an industrial setting and offers a computational advantage that is otherwise impossible to achieve as traditional computational approaches are limited by the choice of approximation scheme employed to discretize the original infinite dimensional behavior. In this way, the paper offers a pathway to exploit new computational techniques in the realm of distributed parameter systems to significantly improve the quality of DoD inkjet printing processes by more accurate modeling and estimation of the thermo-mechanical interactions during fixation.
@@ 2 Preliminaries
Due to confidentiality reasons, in this paper, all the plots are normalized by subtracting the data points with a fixed nominal value. Table 1 shows all the physical constants and their units used in this paper.
Table 1: Physical constants and their units
Symbol Description Units
$\kappa$ Thermal conductivity of the material $W/(m\cdot K)$
$\rho$ Density of the material $kg/m^{3}$
$\chi$ Specific heat capacity of the material $J/(kg\cdot K)$
$h$ Heat transfer coefficient $W/(m^{2}K)$
$g$ Moisture transfer coefficient $m/s$
$\delta$ Moisture diffusion coefficient $m^{2}/s$
The space of functions $f:A\rightarrow B$ is denoted by $(B)^{A}$, i.e., $f\in(B)^{A}$. The array $\begin{bmatrix}a\\ b\end{bmatrix}$ is often written as $\text{col}(a,b)$. $\partial^{q}_{s}f$ denotes the qth-order partial derivative of the function $f$ with respect to $s$, i.e. $\frac{\partial^{q}f}{\partial s^{q}}$.
Given bounded matrix $P\in\mathbb{R}^{n\times m}$, and bounded matrix-valued polynomials $\mathbf{Q_{1}}:[\alpha,\beta]\to\mathbb{R}^{n\times p}$, $\mathbf{Q_{2}}:[\alpha,\beta]\to\mathbb{R}^{q\times m}$, $\mathbf{R_{0}}$$\mathbf{R_{1}}$, and $\mathbf{R_{2}}:[\alpha,\beta]\to\mathbb{R}^{q\times p}$, a Partial Integral (PI) operator denoted by $\mathfrak{\P}\hskip-3.44444pt\left[\footnotesize\begin{array}[]{c|c}P&\mathbf{Q_{1}}\\ \hline\cr\mathbf{Q_{2}}&\{\mathbf{R_{i}}\}\end{array}\right]$ is defined for all $s\in[\alpha,\beta]$, $u\in\mathbb{R}^{m}$ and $\mathbf{v}:[\alpha,\beta]\to\mathbb{R}^{p}$ according to
@@ 3.1 Graph theoretic representation
$\mathcal{G}=(\mathbb{N},A,\mathbb{E}),$ (5)
@@ 3.1.1 Node Structure:
$\mathcal{N}_{i}=(\mathbb{X}_{i},\mathbb{X}_{i}^{\mathrm{bc}},\mathfrak{S}_{i},\mathfrak{P}_{i},\mathfrak{P}_{i}^{\mathrm{bc}},\mathfrak{D}_{i}),$ (6)
@@
$z_{i}$ Estimated output(C) $\mathbb{T}\rightarrow\mathbb{R}^{n_{z,i}}$ Performance index
@@ 3.1.2 Edge Structure:
$\mathcal{E}_{i,j}=(\mathbb{X}_{i,j}^{I},\mathfrak{C}_{i,j}^{I},\mathfrak{E}_{i,j}),$ (7)
@@
Remark 2
Definition 1 provides the blueprint on how to implement the digital twin in software. In fact, following principles of object oriented programming, one may create a node class and edge class with attributes defined according to each item in the definition (6)-(7). For the fixation unit’s model, there are six nodes and each of them are defined using six individual node objects with all the attributes specified. Similar approach is followed for declaring individual edges.
@@ 3.2 Graph-Theoretic Model of the Fixation Unit
Figure 2: Graph theoretic representation of the paper conveyor system. The nodes are color-coded where X, X, and X denote the paper layers, the conveyor belt, and the surrounding air layers, respectively. The arrows show the bi-directional interconnection between the same kind of states The signal $w_{1}(t)$ represents the unknown disturbance acting on the system, while $d_{1}(t)$ and $d_{6}(t)$ are known boundary inputs. The operator $\Delta$ captures the nonlinear evaporation effects, coupling the temperature and moisture dynamics.
The paper sheet is modeled as a non-homogeneous composite with layers $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$, while node $\mathcal{N}_{5}$ represents the conveyor, where only temperature is defined. The dynamics at each spatially distributed node are governed by Fick’s law, subject to interface and external boundary conditions. Thermal contact at interfaces may be either perfect or imperfect. Specifically, the interaction between the paper and the conveyor is modeled with imperfect thermal contact due to energy loss, whereas the thermal contact between the paper layers is assumed to be perfect. External boundaries may also experience disturbances that affect the system.
Remark 3
Only the transversal spatial coordinate is considered because longitudinal variations are significantly smaller during fixation. Moisture transport is dominated by evaporation at the exposed surfaces, making the through-thickness direction the most influential for both drying and thermal transport, while longitudinal moisture gradients remain negligible. Additionally, the conveyor belt has low thermal conductivity and a small cross-sectional area, resulting in minimal temperature variation along its length over the short fixation time. Consequently, longitudinal variations can be neglected for computational efficiency without sacrificing physical accuracy.
@@
which corresponds to the temperature at the bottom surface of the paper. The estimation of $z_{3}(t)$ is chosen because it is the only thermally significant internal quantity that cannot be measured during fixation, yet it provides the missing boundary information needed to reconstruct the full through-thickness temperature profile. This profile is essential for predicting evaporation, moisture migration, and overall drying efficiency. Furthermore, after the first fixation pass, the paper is flipped for the second-side print, at which point the bottom-surface temperature becomes directly measurable. This makes $z_{3}(t)$ a physically verifiable internal variable, enabling validation of both the digital twin and the proposed state-estimation framework.
@@
The moisture content in the printed paper disappears through evaporation in the fixation process. The mass flux of this evaporation depends on both the local paper temperature and moisture content and is captured by the static nonlinear function $\Delta_{1}(p(t))$. Likewise, the associated enthalpy of evaporation depends on the temperature and is represented by the nonlinear term $\Delta_{2}(p(t))$. Although the linear diffusion operators for temperature and moisture are block-diagonal, these nonlinear boundary terms introduce the necessary thermo–fluidic coupling. Temperature influences moisture loss through $\Delta_{1}$, and the resulting latent heat effects feed back into the heat equation through $\Delta_{2}$. Introducing coupling directly into the diffusion matrix would require non-diagonal coefficients with units inconsistent with $m^{2}/s$, rendering such a formulation physically incorrect. Modeling the interaction through $\Delta_{1}$ and $\Delta_{2}$ therefore yields a dimensionally consistent and physically accurate representation of the coupled heat–moisture dynamics.
Based on this coupled thermo–fluidic model, the proposed digital twin reconstructs the missing temperature and moisture profiles inside the paper. The diffusion dynamics propagate the available thermal information across the paper thickness, enabling the internal temperature distribution to be inferred even though it cannot be directly measured during fixation. This reconstructed temperature field is then used in the evaporation relations $\Delta_{1}$ and $\Delta_{2}$ to compute the local moisture loss and the associated latent-heat exchange, resulting in the full spatio-temporal moisture evolution. In this way, the digital twin provides the complete temperature and moisture fields that are otherwise unobservable, effectively functioning as a virtual sensing framework. This is represented by the Linear Fractional Representation (LFR). The exact $\Delta_{1}(p(t))$ and $\Delta_{2}(p(t))$ functions are not mentioned as it is an intellectual property of Canon Production Printing.
Figure 3: Equivalence between the developed model and PIEs.
@@
3.3 Conversion of thermo-fluidic process model into Partial Integral Equations (PIEs)
While the graph-theoretic model as well as the coupled model depicted in Fig. 4 involve boundary conditions including external inputs as well as smoothness conditions in terms of the existence of $\partial_{s}^{2}\mathbf{x}$, as shown in sachin, a wide class of PDE models admits a PIE representation that is free of boundary and smoothness conditions and is parametrized by PI operators. PIE representation and PI operators provide the following advantages that are suitable for the digital twin’s utility:
  • Like matrices, composition, summation, concatenation, and adjoint of PI operators yield a PI operator with analytical formulae for each of these operations. This property is the basis for all algorithmic analysis and synthesis procedures of a given PIE and why PIE framework is advantageous for building the digital twin.
  • In addition to analysis and synthesis, the constraint-free PIE representation also allows for a generalized numerical scheme to approximate the bounded PI operators and simulate PDE models (PEET2024115673) while the original formulation would have required model-specific numerical scheme (e.g., in compared to solving parabolic PDEs, hyperbolic PDEs often require upwind and downwind scheme).
  • All analysis, synthesis, and simulation methods are packaged in an open-source toolbox called PIETOOLS (https://control.asu.edu/pietools/pietools,
    also see 9147712) where the user may define any compatible PDE model which, in turn, gets automatically converted to PIEs and any required analysis, synthesis, and simulation performed as needed on the basis of the PI operator algebra.
  • Finally, under specific conditions, the behavior of the original PDE model is equivalent to the converted PIE model. This means any stability property (PEET2021109473), bound on $L_{2}$ gain between specific input-output transfer ($\mathcal{H}_{\infty}$ norm) (9030224), performance of an optimal observer (das_2019CDC) or a controller (11404177) that are computed in PIETOOLS are equivalently valid for the original PDE model (see sachin for more details).
@@ 3.3 Conversion of thermo-fluidic process model into Partial Integral Equations (PIEs)
Here, $\mathscr{T},\mathscr{T}_{w},\mathscr{T}_{d}$, $\mathscr{A},\mathscr{B}_{11},\mathscr{B}_{12},\mathscr{B}_{q}$, $\mathscr{C}_{1},\mathscr{D}_{11},\mathscr{D}_{12},\mathscr{D}_{1q}$, $\mathscr{C}_{2},\\ \mathscr{D}_{21},\mathscr{D}_{22},\mathscr{D}_{2q}$, $\mathscr{C}_{p},\mathscr{D}_{p1},\mathscr{D}_{p2},\mathscr{D}_{pq}$ are specific PI operators of appropriate dimensions.
@@
The process of constructing the PIEs from the graph-theoretic representation consists of three steps and is depicted in Fig. 3.
@@
For a well-defined PDE, there must be a sufficient number of boundary conditions. Without this, PDE models do not admit a PIE representation. Whether a set of boundary condition is admissible for PIE conversion or not can be easily tested by the invertibility of the matrix $B_{T}$ as given by (phdthesis, Theorem 6.2) where
@@
Since the $B_{T}$ is invertible, using the formulae listed in (phdthesis, Theorm 6.3), the PIE operators $\mathscr{T},\mathscr{T}_{w},\mathscr{T}_{d}$, $\mathscr{A},\mathscr{B}_{11},\mathscr{B}_{12},\mathscr{B}_{q}$, $\mathscr{C}_{1},\mathscr{D}_{11},\mathscr{D}_{12},\mathscr{D}_{1q}$, $\mathscr{C}_{2},\mathscr{D}_{21},\mathscr{D}_{22},\mathscr{D}_{2q}$, $\mathscr{C}_{p}$, $\mathscr{D}_{p1},\mathscr{D}_{p2},\mathscr{D}_{pq}$ are obatined.
@@ 3.4 Implementation and simulation of the model
Based on graph-theoretic definitions, all signals associated with each node are known. The digital twin is implemented in an object-oriented framework using MATLAB and built on top of PIETOOLS version 2022. Each node is represented as a class encapsulating its properties to reduce complexity. The layer class defines each paper layer with parameters such as diffusion, transport, reaction, and transfer coefficients, spatial domain, boundary/initial conditions, and boundary inputs. The spatialconveyor class defines spatial conveyor nodes, storing parameters such as diffusion coefficient, domain size, and initial conditions. This modular design allows flexible definition of any number of states per node in arbitrary order, while ensuring correct interconnections through the $E_{i,j}$ matrix between adjacent nodes. The thermo-fluidic process is simulated using the equivalent PIE formulation from (14), solved via PIE-Galerkin Projection (PGP) (PEET2024115673).